OpenStudy (anonymous):

Where does the graph of y = 6x^2 − 11x − 10 cross the x-axis? (−2/3, 0) and (5/2, 0) (2/3, 0) and (−5/2, 0) (−5/6, 0) and (2, 0) (5/6, 0) and (−2, 0)

OpenStudy (anonymous):

\[x=(-b+or-\sqrt{b^2-4ac})/2a\]

OpenStudy (anonymous):

-b= 11 b^2= 121 a=6 c=-10 2a=12

OpenStudy (anonymous):

\[x= (11 (plus or Minus)\sqrt{121- (4)(6)(-10)})\div12\]

OpenStudy (anonymous):

\[x= (11+-\sqrt{121+240})\div12\]

OpenStudy (anonymous):

\[x= (11+-\sqrt{361})\div12\]

OpenStudy (anonymous):

(11+19)/12 = 5/2 -----> x= 5/2 (11-19)/12 = -2/3 -----> x=-2/3